Superior Company provided the following data for the year ended December 31 (all raw materials are used in production as direct materials): Selling expenses Purchases of raw materials Direct labor 219,000 261,000 Manufacturing overhead applied to work in process Actual manufacturing over s 155,000 $367,000 $355,000 Inventory balances at the beginning and end of the year were as follows: of rear End of Year $ 57,000 37,000 25,000 Work in process Pinished goods $ 38,000 The total manufacturing costs for the year were $680,000; the cost of goods available for sale totaled $745,000: the unadjusted cost of goods sold totaled $665,000; and the net operating income was $32,000. The companys underapplied or overapplied overhead is closed to Cost of Goods Sold. Required Prepare schedules of cost of goods manufactured and cost of goods sold and an income statement (Hint: Prepare the income statement and schedule of cost of goods sold first followed by the schedule of cost of goods manufactured.) Complete this question by entering your answers in the tabs below COGs COGM Prepare an income statement for the year. State
Answer #1
Schedule of Cost of Goods Manufactured
Direct materials:
Raw materials inventory, beginning 57000
Add: Purchases of raw materials 261000
Total raw materials available 318000
Deduct: Raw materials inventory, ending 37000
Raw materials used in production 281000
Direct labor 32000 Balancing Figure =Total Manufacturing Cost-Manufacturing OH applied to inventory-Raw material used in
Manufacturing overhead applied to work in process inventory 367000 production
Total manufacturing costs 680000
Add: Beginning work in process inventory 52000 Balancing Figure
732000 Cost of Goods Manufactured+Ending W.I.P -Total Manufacturing Cost
Deduct: Ending work in process inventory 25000
Cost of goods manufactured 707000
Schedule of Cost of Goods Sold
Finished goods inventory, beginning 38000
Add: Cost of goods manufactured 707000 Balancing Figure
Cost of goods available for sale 745000
Deduct: Finished goods inventory, ending 80000 Balancing Figure
Un adjusted cost of goods sold 665000
Less: Over applied overhead 12000
Adjusted cost of goods sold 653000
Income Statement
Sales 1059000
Cost of goods sold 653000
Gross margin . 406000
Selling and administrative expenses: Sales = Gross margin+Cost of goods sold = 406000+653000=1059000 Gross margin = Net operating income+Selling and Administrative expense = 374000+32000 = 406000 Cost of goods sold = Un adjusted cost of goods sold-Over applied overhead = 665000-12000 = 653000 Over applied overhead = 367000-355000 = $12000
Selling expenses 219000
Administrative expense 155000 374000
Net operating income 32000